I don't know where to start. Can you explain what I am meant to be doing to solve these two questions?
a) 3x + 2y = 5
y = 5x - 17
b) 2x + 3y = 21
x - y = 18

Answer by stanbon(75887)

a) 3x + 2y = 5
y = 5x - 17
-------------
Substitute y=5x-1 into the 1st equation in order to solve for "x":
3x + 2(5x-17) = 5
3x + 10x-34 = 5
13x=39
x=3
--------
Substitute that value of x into y=5x-17 to solve for "y":
y=5*3-17
y=-2
-------
Solution x=3 and y=-2
================================


b) 2x + 3y = 21
x - y = 18
Solve for y: y=x-18
Substitute in the 1st equation to solve for "x":
2x+3(x-18)=21
2x+3x-54=21
5x=75
x=15
Substitute that into y=x-18 to solve for "y":
y=15-18
y=-3
-------
Solution: x=15; y=-3
=========
Cheers,
Stan H.